# On Structural Properties of Risk-Averse Optimal Stopping Problems

### v2

@@ 1 Introduction
In this paper, we address this gap for finite-horizon risk-averse optimal stopping under time-consistent dynamic risk measures [3], focusing on value function monotonicity and the existence of control limit optimal policies. Compared with general MDPs, optimal stopping has a distinct control and cost structure: only two actions (stop or continue), with continuation yields uncontrolled Markovian evolution. In applications, the continuation cost usually reflects only short-term (e.g., one-period) effects, whereas the terminal cost aggregates long-run effects and can be orders of magnitude larger (e.g., [8]); in some financial models [9], only a terminal cost is present. In Section 4, we show that these distinctions make frequently used conditions for identifying structure in general MDPs inapplicable. Although existing studies [22] provide general guidelines tailored to risk-neutral optimal stopping, extending these results to the risk-averse setting is nontrivial because (i) expectation is additive whereas coherent risk measures are only subadditive, and (ii) (conditional) risk measures do not satisfy the tower property—so risk-neutral arguments based on additivity or iterated conditioning break down. Even so, DP remains valid in the risk-averse framework [3], and thus backward induction—the central technique for establishing structural results—is still applicable, making analogous structural results achievable with assumptions and proofs refined for the risk-averse framework. We summarize our main contributions as follows:
@@ 2.1 Optimal Stopping Problem Formulation
Consider a finite time-horizon $\{0,\dots,T\}$ and a Markov process $\{X_{t}\}_{t=0}^{T}$ on a probability space $(\Omega,\mathcal{F},P_{0})$ with reference measure $P_{0}$. Let $\mathcal{X}\cup\{\mathcal{T}\}$ be the state space, where $\mathcal{X}\subseteq{\mathbb{R}}^{n}$ and $\mathcal{T}$ is an absorbing terminal state (no further cost is incurred there). Let $\mathcal{P}(\mathcal{X})$ denote the set of probability measures on $\mathcal{X}$. At each $t\in\{0,\cdots,T-1\}$, the decision-maker chooses an action $u_{t}\in\mathcal{A}=\{S,C\}$ (the action space), where $S$ (stop) terminates the process and moves the state to $\mathcal{T}$, and $C$ (continue) advances the process according to a (possibly time-dependent) Markov transition kernel $Q_{t}(\cdot|x_{t})\in\mathcal{P}(\mathcal{X})$ that depends on the current state $X_{t}=x_{t}\in\mathcal{X}$. Let $\{\mathcal{F}_{t}\}_{t=0}^{T}$ be the filtration adapted to $\{X_{t}\}_{t=0}^{T}$. Given $X_{t}=x_{t}\in\mathcal{X}$, if stopping at time $t$, a terminal cost $s_{t}(x_{t})$ is incurred; otherwise, a continuation cost $c_{t}(x_{t})$ is incurred, where $s_{t},c_{t}:\mathcal{X}\mapsto{\mathbb{R}}$. Define the one-period cost $z_{t}(x,u):=s_{t}(x){\mathbf{1}}\{u=S\}+c_{t}(x){\mathbf{1}}\{u=C\}$. We consider deterministic Markov policies $\mathcal{D}:=\{d=(d_{0},\cdots,d_{T-1})\ |\ d_{t}:\mathcal{X}\mapsto\mathcal{A}\}$. For a policy $d\in\mathcal{D}$, define the stopping time $\tau_{d}:=\min\{t\leq T\ |\ d_{t}(X_{t})=S\}$ with respect to (w.r.t.) $\{\mathcal{F}_{t}\}_{t=0}^{T}$. Let $\mathcal{Z}_{t}$ denote the space of $\mathcal{F}_{t}$-measurable random variables and $Z_{t}:=z_{t}(X_{t},d(X_{t})){\mathbf{1}}\{t\leq\tau\}\in\mathcal{Z}_{t}$ as the period-$t$ cost. The total cost is $\sum_{t=1}^{T}Z_{t}=s_{\tau}(X_{\tau})+\sum_{t=0}^{\tau-1}c_{t}(X_{t})$.
2.2 Time-Consistent Dynamic Risk Measures
@@
Let $\tilde{X}_{t+1}(x)$ denote a random variable with law $Q_{t}(\cdot|x)$, denoted by $\tilde{X}_{t+1}(x)\sim Q_{t}(\cdot|x)$, representing the transition from state $x$ at time $t$. Under suitable regularity assumptions on the transition kernels, one-step risk mappings, and cost functions [3], the risk minimization problem $v_{0}(x):=\min_{d\in\mathcal{D}}J(d,x)$ is solved by the DP recursion
@@ 2.3 Dual Representation of Coherent Risk Measures
For $W\in\mathcal{Z}_{t+1}$ and tail level $\alpha\in(0,1]$, the conditional CVaR at time $t$, denoted by $\operatorname*{CVaR}_{\alpha,t}(W)$, is the conditional mean of the worst $\alpha$-tail of $W$. $\operatorname*{CVaR}_{\alpha,t}$ is coherent and admits the Rockafellar–Uryasev representation [24]: $\operatorname*{CVaR}_{\alpha,t}(W)=\inf_{U\in\mathcal{Z}_{t}}\{U+{\mathbb{E}}((W-U)_{+}|\mathcal{F}_{t})/\alpha\}$. In 3, with $\rho_{t}:=\operatorname*{CVaR}_{\alpha,t}$, the associated risk envelope is $\mathcal{A}_{t}(Q_{t}(\cdot|x_{t}))=\{P\in\mathcal{\mathcal{}}P(\mathcal{X})\ |\ P\ll Q_{t}(\cdot|x_{t}),~dP/dQ_{t}(\cdot|x_{t})\leq 1/\alpha\}$, where $dP/dQ_{t}(\cdot|x_{t})$ is the Radon–Nikodym derivative (the likelihood ratio—the ratio of probability mass functions (pmfs) in the discrete case or of densities in the continuous case), and $P\ll Q_{t}(\cdot|x_{t})$ denotes that $P$ is absolutely continuous w.r.t. $Q_{t}(\cdot|x_{t})$. $\square$
@@ 2.4 Breakdown of Tower Property and Additivity for Coherent Risk Measures
The absence of the tower property and additivity makes some arguments commonly used to establish structural results in risk-neutral settings inapplicable. Without the tower property, one cannot “analyze the target dimension by conditioning on (fixing) the rest,” a standard step for proving structure in a designated dimension (e.g., [22, Proposition 3] and [11, Proposition 1]). Likewise, nonadditivity makes it impossible to apply horizon-wise backward induction to propagate (super/sub)modularity of Q-functions—arguments frequently employed to prove monotone optimal policies [16]. We will revisit these distinctions when proving the specific structural results. In Sections 3 and 4, we show that analogous structural results remain attainable, but require assumptions and/or arguments refined for risk-averse settings.
@@ 3 Monotonicity of the value function
The definition above generalizes the classical notion of FOSD on a totally ordered space [25]: for real-valued random variables $X_{1},X_{2}$, we have $X_{1}\succeq_{sd}X_{2}$ if and only if $F_{1}(x)\leq F_{2}(x)$, or equivalently, $\overline{F}_{1}(x)\geq\overline{F}_{2}(x)$ for all $x\in{\mathbb{R}}$, where $F_{i}$ is the cumulative distribution function (CDF) of $X_{i}$ and $\overline{F}_{i}:=1-F_{i}$. Stochastic monotonicity captures that a smaller current state shifts the conditional next-state distribution toward smaller states and is widely used, e.g., in reliability engineering. As with expectation, coherent risk measures preserve stochastic ordering when they are distribution-invariant (introduced below); commonly used coherent measures—e.g., CVaR, entropic value-at-risk (EVaR), and mean–semideviation—are distribution-invariant.
@@
A one-step risk mapping $\rho_{t}:\mathcal{Z}_{t+1}\mapsto\mathcal{Z}_{t}$ is distribution-invariant if $Z\stackrel{{\scriptstyle d}}{{=}}W$ implies $\rho_{t}(Z)=\rho_{t}(W)$ for all $Z,W\in\mathcal{Z}_{t+1}$, where $\stackrel{{\scriptstyle d}}{{=}}$ denotes “equal in distribution”.
@@ 3.1 Monotonicity in All State Dimensions
For the case where the transition kernel is stochastically monotone on the entire $\mathcal{X}$, the classical risk-neutral argument carries over to the risk-averse setting [13]: if the one-step cost functions are monotone and the kernel is stochastically monotone, then the value function inherits monotonicity. We now state the assumptions and results. For the remainder of the paper, let $\preceq$ denote the componentwise partial order on $\mathcal{X}\subseteq{\mathbb{R}}^{n}$.
@@
Under Assumptions 1 and 2, $v_{t}(x)$ is decreasing in $x$ for all $t$.
@@
We argue by backward induction. Base case: For $t=T$, $v_{T}(x)=s_{T}(x)$ is decreasing by assumption. Inductive step: Suppose $v_{t+1}(x)$ is decreasing for some $t<T$. Fix $x\preceq x^{\prime}$. Since $Q_{t}(\cdot|x)$ is stochastically increasing, we have $\tilde{X}_{t+1}(x)\preceq_{sd}\tilde{X}_{t+1}(x^{\prime})$. By the induction hypothesis, $v_{t+1}$ is decreasing, hence $v_{t+1}\big(\tilde{X}_{t+1}(x)\big)\succeq_{sd}v_{t+1}\big(\tilde{X}_{t+1}(x^{\prime})\big)$. By Lemma 1, $\rho_{t}\big(v_{t+1}\big(\tilde{X}_{t+1}(x)\big)\big)\geq\rho_{t}\big(v_{t+1}\big(\tilde{X}_{t+1}(x^{\prime})\big)\big)$, so $x\mapsto\rho_{t}\big(v_{t+1}\big(\tilde{X}_{t+1}(x)\big)\big)$ is decreasing. Since $c_{t}$ and $s_{t}$ are decreasing, and the minimum of decreasing functions is decreasing, we conclude $v_{t}$ is decreasing.
@@
The risk envelope $\mathcal{A}_{t}(Q_{t}(\cdot|x))$ has a minimal element $\tilde{Q}_{t}(\cdot|x)$ w.r.t. FOSD for each $t$ and $x\in\mathcal{X}$.
@@
By Theorem 1, $v_{t}$ is decreasing. Under Assumption 3, by the definition of FOSD in Definition 2, the supremum in 3 is attained at $\tilde{Q}_{t}(\cdot|x_{t})$ for any decreasing function $v$. Hence, $\rho_{t}\big(v_{t}\big(\tilde{X}_{t+1}(x_{t})\big)\big)={\mathbb{E}}_{\tilde{X}_{t+1}\sim\tilde{Q}_{t}(\cdot|x)}\big(v_{t}\big(\tilde{X}_{t+1}\big)\big).$ Consequently, the DP recursion 2 becomes $v_{t}(x_{t})=\min\{s_{t}(x_{t}),c_{t}(x_{t})+{\mathbb{E}}_{X_{t+1}\sim\tilde{Q}_{t}(\cdot|x_{t})}s_{t+1}(X_{t+1})\}$ for $t<T$ and $v_{T}(x_{T})=s_{T}(x_{T})$, which is exactly the DP recursion of a risk-neutral MDP with transition kernel $\tilde{Q}_{t}(\cdot|x)$ and the same cost functions.
@@ 3.2 Monotonicity in a Subset of Dimensions
$\mathcal{X}=\mathcal{X}_{1}\times\mathcal{X}_{2}$, where $\mathcal{X}_{1}\subseteq{\mathbb{R}}^{n_{1}}$ is equipped with a partial order $\preceq_{1}$, and $\mathcal{X}_{2}\subseteq{\mathbb{R}}^{n_{2}}$, with $n_{1},n_{2}\geq 1$.
@@
Notice that $(\preceq_{1},=)$ defines a partial order on $X_{1}\times\mathcal{X}_{2}$; see [11].
@@
  • Given $x_{2}$ and conditioning on $\tilde{X}_{t,2}(x)$, $\tilde{X}_{t,1}(x_{1},x_{2})$ is stochastically increasing in $x_{1}$, i.e., $x_{1}\preceq_{1}x_{1}^{\prime}\implies\tilde{X}_{t,1}(x_{1},x_{2})\preceq_{sd}\tilde{X}_{t,1}(x_{1}^{\prime},x_{2})$.
@@
Under Assumptions 1, 4, and 5, $v_{t}(x_{1},x_{2})$ is decreasing in $x_{1}$ for all $t$.
@@
We argue by backward induction on $t$. Base case: For $t=T$, $v_{T}(x_{1},x_{2})=s_{T}(x_{1},x_{2})$ is decreasing in $x_{1}$ by assumption. Inductive step: Fix $t<T$ and assume $v_{t+1}(x_{1},x_{2})$ is decreasing in $x_{1}$ for every $x_{2}$. Let $x=(x_{1},x_{2})$ and $x^{\prime}=(x_{1}^{\prime},x_{2})$ with $x_{1}\preceq_{1}x_{1}^{\prime}$. By Assumption 5, we have $\tilde{X}_{t+1,2}(x)=\tilde{X}_{t+1,2}(x^{\prime})$ a.s., and $\tilde{X}_{t+1,1}(x)\preceq_{sd}\tilde{X}_{t+1,1}(x^{\prime})$. Construct coupled next-state vectors $\hat{X}_{t+1}(x)=\big(\hat{X}_{t+1,1}(x),\hat{X}_{t+1,2}(x_{2})\big)$, $\hat{X}_{t+1}(x^{\prime})=\big(\hat{X}_{t+1,1}(x^{\prime}),\hat{X}_{t+1,2}(x_{2})\big)$ as follows: (i) Synchronize the second coordinate by taking $\hat{X}_{t+1,2}(x_{2})=\tilde{X}_{t+1,2}(x_{2})$ a.s.; (ii) Conditioning on $\hat{X}_{t+1,2}(x_{2})$, consider the first coordinate. Under Assumption 5 (third bullet), by Strassen’s theorem [27, 28], there exists random variables $\hat{X}_{t+1,1}(x)$ and $\hat{X}_{t+1,1}(x^{\prime})$ such that (i) $\hat{X}_{t+1,1}(x)\stackrel{{\scriptstyle d}}{{=}}\tilde{X}_{t+1,1}(x)$, (ii) $\hat{X}_{t+1,1}(x^{\prime})\stackrel{{\scriptstyle d}}{{=}}\tilde{X}_{t+1,1}(x^{\prime})$, and (iii) $\hat{X}_{t+1,1}(x)\preceq\hat{X}_{t+1,1}(x^{\prime})$ a.s. Consequently, $\hat{X}_{t+1}(x)\stackrel{{\scriptstyle d}}{{=}}\tilde{X}_{t+1}(x)$, $\hat{X}_{t+1}(x^{\prime})\stackrel{{\scriptstyle d}}{{=}}\tilde{X}_{t+1}(x^{\prime})$, and $\hat{X}_{t+1,2}(x)=\hat{X}_{t+1,2}(x^{\prime})$, $\hat{X}_{t+1,1}(x)\preceq_{sd}\hat{X}_{t+1,1}(x^{\prime})$ a.s. Since $v_{t+1}(x_{1},x_{2})$ is decreasing in $x_{1}$, we have $v_{t+1}\big(\hat{X}_{t+1}(x)\big)\geq v_{t+1}\big(\hat{X}_{t+1}(x^{\prime})\big)$ a.s., and thus $\rho_{t}(v_{t}(\hat{X}_{t}(x)))\geq\rho_{t}(v_{t}(\hat{X}_{t+1}(x^{\prime})))$. Since $\rho_{t}$ is distribution-invariant, by Lemma 1, $\rho_{t}\big(v_{t+1}(\tilde{X}_{t+1}(x))\big)\geq\rho_{t}\big(v_{t+1}(\tilde{X}_{t+1}(x^{\prime}))\big)$, so $\rho_{t}\big(v_{t+1}(\tilde{X}_{t+1}(x))\big)$ is decreasing in $x_{1}$. Finally, since $s_{t}(\cdot,x_{2})$ and $c_{t}(\cdot,x_{2})$ are decreasing for fixed $x_{2}$ and the minimum of decreasing functions is decreasing, the DP recursion (2) implies that $v_{t}(x_{1},x_{2})$ is decreasing in $x_{1}$.
The following average-rate forward (ARF) model with an early-termination feature illustrates Theorem 3.7. The Asian–American call option has the same two-dimensional state dynamics but a slightly different option-style terminal payoff [9]; the analysis and results below carry over.
@@
Indeed, in this example, $v_{t}(x)$ is decreasing in both coordinate; this also follows from Theorem 1, since $\tilde{X}_{t+1}(x_{t})$ is stochastically increasing in $x_{t}$ under the componentwise partial order on ${\mathbb{R}}^{2}$. $\square$
For another example satisfying Assumption 5, see the organ transplantation model in [8], where $X_{t}=(X_{t,1},X_{t,2})$: $X_{t,2}$ is the patient’s health state, evolving only from the previous health, and $X_{t,1}$ is the organ-offer quality state, which is stochastically monotone in health state (patients in better health are more likely to receive higher-quality offers).
4 Characterize the structure of the optimal policy
In this section, we establish conditions for the existence of control limit optimal policies, beginning with a precise definition. Let $x=(x_{1},\cdots,x_{n})\in\mathcal{X}\subseteq{\mathbb{R}}^{n}$ and write $x_{-i}$ for the vector with its $i$-th component removed. Consider the componentwise partial order $\preceq$ on $\mathcal{X}\subseteq{\mathbb{R}}^{n}$.

### v3

@@ 1 Introduction
In this paper, we address this gap for finite-horizon risk-averse optimal stopping under time-consistent dynamic risk measures [3], focusing on value function monotonicity and the existence of control limit optimal policies. Compared with general MDPs, optimal stopping has a distinct control and cost structure: only two actions (stop or continue), with continuation yielding uncontrolled Markovian evolution. In applications, the continuation cost usually reflects only short-term (e.g., one-period) effects, whereas the terminal cost aggregates long-run effects and can be orders of magnitude larger (e.g., [8]); in some financial models [9], only a terminal cost is present. In Section 4, we show that these distinctions make frequently used conditions for identifying structure in general MDPs inapplicable. Although existing studies [22] provide general guidelines tailored to risk-neutral optimal stopping, extending these results to the risk-averse setting is nontrivial because (i) expectation is additive whereas coherent risk measures are only subadditive, and (ii) (conditional) risk measures do not satisfy the tower property—so risk-neutral arguments based on additivity or iterated conditioning break down. Even so, DP remains valid in the risk-averse framework [3], and thus backward induction—the central technique for establishing structural results—is still applicable, making analogous structural results achievable with assumptions and proofs refined for the risk-averse framework. We summarize our main contributions as follows:
@@ 2.1 Optimal Stopping Problem Formulation
Consider a finite time-horizon $\{0,\dots,T\}$ and a Markov process $\{X_{t}\}_{t=0}^{T}$ on a probability space $(\Omega,\mathcal{F},P_{0})$ with reference measure $P_{0}$. Let $\mathcal{X}\cup\{\mathcal{T}\}$ be the state space, where $\mathcal{X}\subseteq{\mathbb{R}}^{n}$ and $\mathcal{T}$ is an absorbing terminal state (no further cost is incurred there). Let $\mathcal{P}(\mathcal{X})$ denote the set of probability measures on $\mathcal{X}$. At each $t\in\{0,\cdots,T-1\}$, the decision-maker chooses an action $u_{t}\in\mathcal{A}=\{S,C\}$ (the action space), where $S$ (stop) terminates the process and moves the state to $\mathcal{T}$, and $C$ (continue) advances the process according to a (possibly time-dependent) Markov transition kernel $Q_{t}(\cdot|x_{t})\in\mathcal{P}(\mathcal{X})$ that depends on the current state $X_{t}=x_{t}\in\mathcal{X}$. Let $\{\mathcal{F}_{t}\}_{t=0}^{T}$ be the filtration generated by $\{X_{t}\}_{t=0}^{T}$. Given $X_{t}=x_{t}\in\mathcal{X}$, if stopping at time $t$, a terminal cost $s_{t}(x_{t})$ is incurred; otherwise, a continuation cost $c_{t}(x_{t})$ is incurred, where $s_{t},c_{t}:\mathcal{X}\mapsto{\mathbb{R}}$. Define the one-period cost $z_{t}(x,u):=s_{t}(x){\mathbf{1}}\{u=S\}+c_{t}(x){\mathbf{1}}\{u=C\}$. We consider deterministic Markov policies $\mathcal{D}:=\{d=(d_{0},\cdots,d_{T-1})\ |\ d_{t}:\mathcal{X}\mapsto\mathcal{A}\}$. For a policy $d\in\mathcal{D}$, define the stopping time $\tau_{d}:=\min\{t\leq T\ |\ d_{t}(X_{t})=S\}$ adapted to $\{\mathcal{F}_{t}\}_{t=0}^{T}$. Let $\mathcal{Z}_{t}$ denote the space of $\mathcal{F}_{t}$-measurable random variables and $Z_{t}:=z_{t}(X_{t},d(X_{t})){\mathbf{1}}\{t\leq\tau\}\in\mathcal{Z}_{t}$ as the period-$t$ cost. The total cost is $\sum_{t=1}^{T}Z_{t}=s_{\tau}(X_{\tau})+\sum_{t=0}^{\tau-1}c_{t}(X_{t})$.
2.2 Risk-Averse Optimal Stopping
@@
For fixed $x\in\mathcal{X}$, let $\tilde{X}_{t+1}(x)$ denote a random variable with law $Q_{t}(\cdot|x)$, denoted by $\tilde{X}_{t+1}(x)\sim Q_{t}(\cdot|x)$, representing the transition from state $x$ at time $t$. Under suitable regularity assumptions on the transition kernels, one-step risk mappings, and cost functions [3], the risk minimization problem $v_{0}(x):=\min_{d\in\mathcal{D}}J(d,x)$ is solved by the DP recursion
@@ 2.3 Dual Representation of Coherent Risk Measures
For $W\in\mathcal{Z}_{t+1}$ and tail level $\alpha\in(0,1]$, the conditional CVaR at time $t$, denoted by $\operatorname*{CVaR}_{\alpha,t}(W)$, is the $\mathcal{F}_{t}$-conditional mean of $W$ over its worst $\alpha$-tail. $\operatorname*{CVaR}_{\alpha,t}$ is coherent and admits the Rockafellar–Uryasev representation [24]: $\operatorname*{CVaR}_{\alpha,t}(W)=\inf_{U\in\mathcal{Z}_{t}}\{U+{\mathbb{E}}((W-U)_{+}|\mathcal{F}_{t})/\alpha\}$. In 3, with $\rho_{t}:=\operatorname*{CVaR}_{\alpha,t}$, the associated risk envelope is $\mathcal{A}_{t}(Q_{t}(\cdot|x_{t}))=\{P\in\mathcal{\mathcal{}}P(\mathcal{X})\ |\ P\ll Q_{t}(\cdot|x_{t}),~dP/dQ_{t}(\cdot|x_{t})\leq 1/\alpha\}$, where $dP/dQ_{t}(\cdot|x_{t})$ is the Radon–Nikodym derivative (the likelihood ratio—the ratio of probability mass functions (pmfs) in the discrete case or of densities in the continuous case), and $P\ll Q_{t}(\cdot|x_{t})$ denotes that $P$ is absolutely continuous with respect to (w.r.t.) $Q_{t}(\cdot|x_{t})$. $\square$
@@ 2.4 Breakdown of Tower Property and Additivity for Coherent Risk Measures
The absence of the tower property and additivity makes some arguments commonly used to establish structural results in risk-neutral settings inapplicable. Without the tower property, one cannot “analyze the target dimension by conditioning on (fixing) the rest,” a standard step for proving structure in a designated dimension (e.g., [22, Proposition 3] and [11, Proposition 1]). Likewise, nonadditivity makes it impossible to apply horizon-wise backward induction to propagate (super/sub)modularity of Q-functions—arguments frequently employed to prove monotone optimal policies [16]. We will revisit these distinctions when proving the specific structural results. In Sections 3 and 4, we show that analogous structural results remain attainable, but require assumptions and/or arguments tailored for risk-averse settings.
@@ 3 Monotonicity of the value function
The definition above generalizes the classical notion of FOSD on a totally ordered space [25]: for real-valued random variables $X_{1},X_{2}$, we have $X_{1}\succeq_{sd}X_{2}$ if and only if $F_{1}(x)\leq F_{2}(x)$, or equivalently, $\overline{F}_{1}(x)\geq\overline{F}_{2}(x)$, $\forall x\in{\mathbb{R}}$, where $F_{i}$ is the cumulative distribution function (CDF) of $X_{i}$ and $\overline{F}_{i}:=1-F_{i}$. Stochastic monotonicity captures that a smaller current state shifts the conditional next-state distribution toward smaller states and is widely used, e.g., in reliability engineering. As with expectation, coherent risk measures preserve stochastic ordering when they are distribution-invariant (introduced below); commonly used coherent measures—e.g., CVaR, entropic value-at-risk (EVaR), and mean–semideviation—are distribution-invariant.
@@
A one-step risk mapping $\rho_{t}:\mathcal{Z}_{t+1}\mapsto\mathcal{Z}_{t}$ is distribution-invariant if $Z\stackrel{{\scriptstyle d}}{{=}}W$ implies $\rho_{t}(Z)=\rho_{t}(W),~\forall Z,W\in\mathcal{Z}_{t+1}$, where $\stackrel{{\scriptstyle d}}{{=}}$ denotes “equal in distribution”.
@@ 3.1 Monotonicity in All State Dimensions
For the case where the transition kernel is stochastically monotone on the entire $\mathcal{X}$, the classical risk-neutral argument carries over to the risk-averse setting [13]: if the one-step cost functions are monotone and the kernel is stochastically monotone, then the value function inherits monotonicity. We now state the assumptions and results. Let $\preceq_{c}$ denote the componentwise partial order on $\mathcal{X}\subseteq{\mathbb{R}}^{n}$.
@@
Under Assumptions 1 and 2, $v_{t}(x)$ is decreasing in $x$, $\forall t$.
@@
We argue by backward induction. Base case: For $t=T$, $v_{T}(x)=s_{T}(x)$ is decreasing by assumption. Inductive step: Suppose $v_{t+1}(x)$ is decreasing for some $t<T$. Fix $x\preceq_{c}x^{\prime}$. Since $Q_{t}(\cdot|x)$ is stochastically increasing, we have $\tilde{X}_{t+1}(x)\preceq_{sd}\tilde{X}_{t+1}(x^{\prime})$. By the induction hypothesis, $v_{t+1}$ is decreasing, hence $v_{t+1}\big(\tilde{X}_{t+1}(x)\big)\succeq_{sd}v_{t+1}\big(\tilde{X}_{t+1}(x^{\prime})\big)$. By Lemma 1, $\rho_{t}\big(v_{t+1}\big(\tilde{X}_{t+1}(x)\big)\big)\geq\rho_{t}\big(v_{t+1}\big(\tilde{X}_{t+1}(x^{\prime})\big)\big)$, so $x\mapsto\rho_{t}\big(v_{t+1}\big(\tilde{X}_{t+1}(x)\big)\big)$ is decreasing. Since $c_{t}$ and $s_{t}$ are decreasing, and the minimum of decreasing functions is decreasing, we conclude $v_{t}$ is decreasing.
@@
The risk envelope $\mathcal{A}_{t}(Q_{t}(\cdot|x))$ has a minimal element $\tilde{Q}_{t}(\cdot|x)$ w.r.t. FOSD, $\forall t$ and $x\in\mathcal{X}$.
@@
By Theorem 1, $v_{t}$ is decreasing. Under Assumption 3, by the definition of FOSD in Definition 2, the supremum in 3 is attained at $\tilde{Q}_{t}(\cdot|x_{t})$ for any decreasing function $v$. Hence, $\rho_{t}\big(v_{t}\big(\tilde{X}_{t+1}(x_{t})\big)\big)={\mathbb{E}}_{\tilde{X}_{t+1}\sim\tilde{Q}_{t}(\cdot|x)}\big(v_{t}\big(\tilde{X}_{t+1}\big)\big).$ Consequently, the DP recursion 2 becomes $v_{t}(x_{t})=\min\big\{s_{t}(x_{t}),c_{t}(x_{t})+{\mathbb{E}}_{\tilde{X}_{t+1}\sim\tilde{Q}_{t}(\cdot|x)}\big(v_{t}\big(\tilde{X}_{t+1}\big)\big)\big\}$ for $t<T$ and $v_{T}(x_{T})=s_{T}(x_{T})$, which is exactly the DP recursion of a risk-neutral MDP with transition kernel $\tilde{Q}_{t}(\cdot|x)$ and the same cost functions.
@@ 3.2 Monotonicity in a Subset of Dimensions
$\mathcal{X}=\mathcal{X}_{1}\times\mathcal{X}_{2}$, where $\mathcal{X}_{1}\subseteq{\mathbb{R}}^{n_{1}}$ is equipped with a partial order $\preceq_{1}$, and $\mathcal{X}_{2}\subseteq{\mathbb{R}}^{n_{2}}$, with $n_{1}+n_{2}=n,~n_{1},n_{2}\geq 1,$.
@@
Notice that $(\preceq_{1},=)$ defines a partial order on $\mathcal{X}_{1}\times\mathcal{X}_{2}$; see [11].
Throughout the paper, we say that a multivariable function is monotone in some component if the function is monotone in that component with other components fixed.
@@
  • For fixed $x_{2}$ and conditioning on $\tilde{X}_{t,2}(x)$, $\tilde{X}_{t,1}(x_{1},x_{2})$ is stochastically increasing in $x_{1}$, i.e., $x_{1}\preceq_{1}x_{1}^{\prime}\implies\tilde{X}_{t,1}(x_{1},x_{2})\preceq_{sd}\tilde{X}_{t,1}(x_{1}^{\prime},x_{2})$.
@@
Under Assumptions 1, 4, and 5, $v_{t}(x_{1},x_{2})$ is decreasing in $x_{1}$, $\forall t$.
@@
We argue by backward induction on $t$. Base case: For $t=T$, $v_{T}(x_{1},x_{2})=s_{T}(x_{1},x_{2})$ is decreasing in $x_{1}$ by assumption. Inductive step: Fix $t<T$ and assume $v_{t+1}(x_{1},x_{2})$ is decreasing in $x_{1}$ for each fixed $x_{2}$. Let $x=(x_{1},x_{2})$ and $x^{\prime}=(x_{1}^{\prime},x_{2})$ with $x_{1}\preceq_{1}x_{1}^{\prime}$. By Assumption 5, we have $\tilde{X}_{t+1,2}(x)=\tilde{X}_{t+1,2}(x^{\prime})$ a.s., and $\tilde{X}_{t+1,1}(x)\preceq_{sd}\tilde{X}_{t+1,1}(x^{\prime})$. Construct next-state vectors $\hat{X}_{t+1}(x)=\big(\hat{X}_{t+1,1}(x),\hat{X}_{t+1,2}(x_{2})\big)$, $\hat{X}_{t+1}(x^{\prime})=\big(\hat{X}_{t+1,1}(x^{\prime}),\hat{X}_{t+1,2}(x_{2})\big)$ as follows: (i) Synchronize the second coordinate by taking $\hat{X}_{t+1,2}(x_{2})=\tilde{X}_{t+1,2}(x_{2})$ a.s.; (ii) Conditioning on $\hat{X}_{t+1,2}(x_{2})$, consider the first coordinate. Under Assumption 5 (third bullet), by Strassen’s theorem [27, 28], there exists random variables $\hat{X}_{t+1,1}(x)$ and $\hat{X}_{t+1,1}(x^{\prime})$ such that (i) $\hat{X}_{t+1,1}(x)\stackrel{{\scriptstyle d}}{{=}}\tilde{X}_{t+1,1}(x)$, (ii) $\hat{X}_{t+1,1}(x^{\prime})\stackrel{{\scriptstyle d}}{{=}}\tilde{X}_{t+1,1}(x^{\prime})$, and (iii) $\hat{X}_{t+1,1}(x)\preceq_{1}\hat{X}_{t+1,1}(x^{\prime})$ a.s. Consequently, $\hat{X}_{t+1}(x)\stackrel{{\scriptstyle d}}{{=}}\tilde{X}_{t+1}(x)$, $\hat{X}_{t+1}(x^{\prime})\stackrel{{\scriptstyle d}}{{=}}\tilde{X}_{t+1}(x^{\prime})$, and $\hat{X}_{t+1,2}(x)=\hat{X}_{t+1,2}(x^{\prime})$, $\hat{X}_{t+1,1}(x)\preceq_{sd}\hat{X}_{t+1,1}(x^{\prime})$ a.s. Since $v_{t+1}(x_{1},x_{2})$ is decreasing in $x_{1}$, we have $v_{t+1}\big(\hat{X}_{t+1}(x)\big)\geq v_{t+1}\big(\hat{X}_{t+1}(x^{\prime})\big)$ a.s., and thus $\rho_{t}(v_{t}(\hat{X}_{t}(x)))\geq\rho_{t}(v_{t}(\hat{X}_{t+1}(x^{\prime})))$. Since $\rho_{t}$ is distribution-invariant, by Lemma 1, $\rho_{t}\big(v_{t+1}(\tilde{X}_{t+1}(x))\big)\geq\rho_{t}\big(v_{t+1}(\tilde{X}_{t+1}(x^{\prime}))\big)$, so $\rho_{t}\big(v_{t+1}(\tilde{X}_{t+1}(x))\big)$ is decreasing in $x_{1}$. Finally, since $s_{t}(\cdot,x_{2})$ and $c_{t}(\cdot,x_{2})$ are decreasing for fixed $x_{2}$ and the minimum of decreasing functions is decreasing, the DP recursion (2) implies that $v_{t}(x_{1},x_{2})$ is decreasing in $x_{1}$.
The following average-rate forward (ARF) model with an early-termination feature illustrates Theorem 3.7. The Asian–American call option has the same two-dimensional state dynamics as Example 3.9 but a slightly different option-style terminal payoff [9]; the analysis and results below carry over.
@@
Indeed, in this example, $v_{t}(x)$ is decreasing in each of the coordinates; this also follows from Theorem 1, since $\tilde{X}_{t+1}(x_{t})$ is stochastically increasing in $x_{t}$ under the componentwise partial order on ${\mathbb{R}}^{2}$. $\square$
For another example satisfying Assumption 5, see the organ transplantation model in [8], where $X_{t}=(X_{t,1},X_{t,2})$: $X_{t,2}$ is the patient’s health state, evolving only from the previous health state, and $X_{t,1}$ is the organ-offer quality state, which is stochastically monotone in the health state (patients in better health are more likely to receive higher-quality offers).
4 Characterization of the structure of the optimal policy
In this section, we establish conditions for the existence of control limit optimal policies, beginning with a precise definition. Let $x=(x_{1},\cdots,x_{n})\in\mathcal{X}\subseteq{\mathbb{R}}^{n}$ and define the (n1)(n-1)-dimensional vector $x_{-i}$ to be the vector $x$ with its $i$-th component removed. Throughout this section, we consider monotonicity w.r.t. the componentwise partial order $\preceq_{c}$ on $\mathcal{X}\subseteq{\mathbb{R}}^{n}$.